OpenStudy (anonymous):

Solve the equation for the letter y: -2x + 3y = 4 Please help?

ganeshie8 (ganeshie8):

start by adding 2x both sides

ganeshie8 (ganeshie8):

-2x + 3y = 4 +2x +2x

ganeshie8 (ganeshie8):

wat do u get ? :)

OpenStudy (anonymous):

so it will be 3y=6x ?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

you cant add apples and cocounts

ganeshie8 (ganeshie8):

-2x + 3y = 4 +2x +2x it wud be :- 3y = 4+2x

ganeshie8 (ganeshie8):

you cant add 4 and 2x ok they're not like terms

ganeshie8 (ganeshie8):

next divide 3 both sides, to get y all by itself :- 3y = 4+2x /3 /3 y = (4+2x)/3

ganeshie8 (ganeshie8):

see if that makes more/less sense :)

OpenStudy (anonymous):

so y=2

ganeshie8 (ganeshie8):

nopes, y = (4+2x)/3

ganeshie8 (ganeshie8):

we're done.

ganeshie8 (ganeshie8):